OpenStudy (anonymous):

Please help me simplify... I'll be back in 40 min

OpenStudy (anonymous):

\[3(a^2 b) ^ -3 / (3ab)^2 \] i forgot how to write the -3 power correctly.

OpenStudy (anonymous):

\[\frac{ 3(a^2b)^-3 }{ (3ab)^2 }\] that's a little clearer. :)

OpenStudy (anonymous):

ok ill try one more time to use this equation editor... \[\frac{ 3(a^2b)^{-3} }{(3ab)^2 } \] finally, now could someone simplify this please c?

OpenStudy (anonymous):

I can type out my own working if you reakllyh want me to ?

OpenStudy (anonymous):

is the bottom \[9a^2b^2\]

OpenStudy (anonymous):

You have two rules here (1) When you take an exponent of an exponent they multiple \[{(A)^2}^{5} = (A)^{2*5}\] (2) When you have an exponent being divided by another exponent they subtract \[\frac{A^8}{A^4} = A^{8-4}\]\ Use these rules and match the numbers with the same base

OpenStudy (anonymous):

thank you. grat answer.

OpenStudy (anonymous):

Look : we have : (a^2b)^-2 = a^2*(-2) * b ^-2 = a^-4 * b^-2 3( a^-4 * b^-2 ) = 3a^-4 * b ^ -2 got it ?

OpenStudy (anonymous):

then : ( 3ab)^2 = 3^2 * a^2 * b ^2 = 9a^2B^2 have : 3a^-4 * b ^ -2/9a^2B^2 = 3a^-4/9a^2 * b^-2/b^2 = a^-4/3a^2 * b^-4 = 1/3 * a^-6 * b^-4 Got it ?;)

OpenStudy (anonymous):

i GOT \[\frac{ 3(a^{-6}*b ^{-3}) }{ 9a^2b^2 }\]= \[\frac{ 1*a ^{-6-2}b ^{-3-2} }{ 3 } \] = \[\frac{ 1 }{ 3a^8b^5 }\] so I may have mucked it up, but i think I followed your rules ?